OpenStudy (anonymous):

Let f(x) = -2x - 7 and g(x) = -4x + 3. Find (f o g)(-5) A. 23 B. 3 C. -9 D. -53

OpenStudy (anonymous):

You can do this 2 ways. Let x = -5, and solve for g, and put that number into f(x), or: solve for f(x), letting x = g(x), then substituting x for -5

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

write down the equation for f(x), but replace x with the equation of g(x). f(x) = -2(g(x))-7 f(x) = -2(-4x+3)-7 simplify it, and then solve, letting x = -5

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

B

OpenStudy (anonymous):

@goformit100